OpenStudy (anonymous):

2/3+5/6 2/15+3/5 7/10+3/20 11/15+4/45

OpenStudy (anonymous):

last questions

OpenStudy (anonymous):

2/3 turns into 4/6 4/6+5/6 = 9/6 which converts to 3/2

OpenStudy (anonymous):

3/5 turns into 9/15 2/15+9/15 = 11/15

OpenStudy (anonymous):

7/10 turns into 14/20 14/20+3/20 = 17/20 11/15 turns into 33/45 33/45+4/45 = 37/45